OpenStudy (anonymous):

Will fan & medal Simplify: 2 (^3sqrt108)

OpenStudy (anonymous):

@dan815 @zepdrix @Directrix

Directrix (directrix):

Is this supposed to be: 2 ^ (3sqrt108) ? Two raised to the (3 times square root of 108) power?

OpenStudy (anonymous):

No, its just the number 2 being multiplied by the sqrt of 108 that has ^3 in front of the sqrt sign... I learned in my lesson that it means 108^1/3

OpenStudy (anonymous):

@Directrix

OpenStudy (johnweldon1993):

Oh okay \(\large 2(\sqrt[3]{108})\)

OpenStudy (anonymous):

Yes!!

OpenStudy (johnweldon1993):

So try and think about cubed roots \(\large 2^3 = 8\) <---can this be evenly divided into 108? \(\large 3^3 = 27\) <---what about this? \(\large 4^3 = 64\) <--- or this? We know 64 cannot be evenly divided into 108..so ignore that...but in what way ...can we manipulate the numbers 3 and 2 to break down 108? *Remembering the key is cubed root*

OpenStudy (anonymous):

Hmmm... I'm thrown off by the 3 and 2 you gave me to work with.. Not really sure what I'm supposed to do :/

OpenStudy (johnweldon1993):

Well the reason why I'm giving you that is because if you think about it \[\large \sqrt[3]{x^3} = x\] right? the 3 and the 3 cancel leaving just 'x'...so if we can find something that is raised to the third power....this can be simplified greatly So hint* (kinda a big one lol) \(\large 3^3 * 4\) what does that equal?

OpenStudy (anonymous):

27 times 4 is 108 ;p

OpenStudy (anonymous):

Wait!!

OpenStudy (anonymous):

You may be reading too much into this because the goal is not to cancel out that exponent.. It still remains in the answer

OpenStudy (johnweldon1993):

Exactly :P So if we take your original question \[\large 2(\sqrt[3]{108})\] and write it as \(\large 2\sqrt[3]{(3^3 * 4)}\) that is the same ***and the exponent will still be in the answer btw :P *** just a way of breaking it down

OpenStudy (anonymous):

My answer choices are: 16 3^sqrt18 8 3^sqrt6 5 3^sqrt4 6 3^sqrt4

OpenStudy (anonymous):

Ohh okay haha got it

OpenStudy (johnweldon1993):

mmhmm i see the one i'm aiming at ;P

OpenStudy (anonymous):

Lol!!!

OpenStudy (johnweldon1993):

Okay SO remember how i said \[\large \sqrt[3]{x^3} = x\] Well...lookey what we have here... \[\large 2(\sqrt[\color \red{3}]{\color \red{3^3} * 4})\]

OpenStudy (anonymous):

oh!!! i know the answer now ;p

OpenStudy (johnweldon1993):

So now if we re-write this as \[\large 2(3\sqrt[3]{4})\] I believe you can do the rest ;P

OpenStudy (anonymous):

so its 6 3^sqrt4

OpenStudy (johnweldon1993):

BAM! lol "ju got et chicha" ;p

OpenStudy (anonymous):

Bahaha thanks bby ;*

OpenStudy (johnweldon1993):

Anytime puddin ;* lol